### v1

# Beyond Joint Demonstrations: Personalized Expert Guidance for Efficient Multi-Agent Reinforcement Learning

Abstract
Multi-Agent Reinforcement Learning (MARL) algorithms face the challenge of efficient exploration due to the exponential increase in the size of the joint state-action space. While demonstration-guided learning has proven beneficial in single-agent settings, its direct applicability to MARL is hindered by the practical difficulty of obtaining joint expert demonstrations. In this work, we introduce a novel concept of personalized expert demonstrations, tailored for each individual agent or, more broadly, each individual type of agent within a heterogeneous team. These demonstrations solely pertain to single-agent behaviors and how each agent can achieve personal goals without encompassing any cooperative elements, thus naively imitating them will not achieve cooperation due to potential conflicts. To this end, we propose an approach that selectively utilizes personalized expert demonstrations as guidance and allows agents to learn to cooperate, namely personalized expert-guided MARL (PegMARL). This algorithm utilizes two discriminators: the first provides incentives based on the alignment of policy behavior with demonstrations, and the second regulates incentives based on whether the behavior leads to the desired objective. We evaluate PegMARL using personalized demonstrations in both discrete and continuous environments. The results demonstrate that PegMARL learns near-optimal policies even when provided with suboptimal demonstrations, and outperforms state-of-the-art MARL algorithms in solving coordinated tasks. We also showcase PegMARL’s capability to leverage joint demonstrations in the StarCraft scenario and converge effectively even with demonstrations from non-co-trained policies.
1 Introduction
The use of expert demonstrations has been proven effective in accelerating learning in single-agent reinforcement learning, as evidenced by studies such as Kang et al. (2018); Chen & Xu (2022); Rengarajan et al. (2022). This approach has since been extended to Multi-Agent Reinforcement Learning (MARL) (Lee & Lee, 2019; Qiu et al., 2022), which typically assumes the availability of high-quality collaborative joint demonstrations. However, from a practical standpoint, collecting joint demonstrations can be labor-intensive, demanding one user per agent in cooperative scenarios. Furthermore, these demonstrations are not scalable. If we change the number of agents or introduce new types of agents, we will need to gather a new set of demonstrations to learn from.
In contrast, it is much easier to obtain demonstrations for individual agents, or even better, for each type of agents in a heterogeneous setting. We thus ask the following research question: could we leverage individual-wise task demonstrations instead? In this work, we refer to such expert demonstrations that address single-agent behaviors for personal objectives as personalized demonstrations (Figure 1). Since the personalized demonstrations will not necessarily reflect how the agents can collaborate and may even conflict with each other in the joint setting, naively mimicking the demonstrations will not achieve cooperation. Therefore, purely imitation learning-based approaches would not be effective. We need an approach that selectively utilizes suitable personalized expert demonstrations as guidance and allows agents to learn to cooperate via collecting reward signals from the environments.
Figure 1: Joint demonstrations are costly to collect but offer rich information on collaborative behaviors. Personalized demonstrations are easier to collect, but solely focus on individual agent goals, so they lack cooperative elements.
To this end, we present our algorithm, Personalized Expert-Guided MARL (PegMARL), which carries out reward-shaping as a form of guidance. We implement this via two discriminators. The first, a personalized behavior discriminator, evaluates local state-action pairs, providing positive incentives for actions that align with the demonstration and negative incentives for divergent ones. The second, a personalized dynamics discriminator, assesses whether a local state-action pair induces a desired change in dynamics similar to that observed in the demonstration, adjusting the incentive weight accordingly. We prove the effectiveness of PegMARL on both discrete gridworld and continuous multi-agent particle environments (Lowe et al., 2017; Mordatch & Abbeel, 2017). The main contributions of this paper are as follows:
(1) Our approach is the first to enable utilizing personalized demonstrations for policy learning in heterogeneous MARL environments, regardless of the number and type of agents involved.
(2) We propose PegMARL that dynamically and selectively reshapes the original reward to aid exploration. It is worth mentioning that this algorithm is general and compatible with most MARL policy gradient methods.
(3) We evaluate PegMARL with personalized demonstrations in both discrete and continuous environments. Our algorithm outperforms state-of-the-art decentralized MARL algorithms, pure multi-agent imitation learning, and reward-shaping techniques in terms of scalability and convergence speed, and obtains near-optimal policies even when provided with suboptimal demonstrations.
(4) We showcase PegMARL’s capability to also leverage joint demonstrations, regardless of whether they are sampled from co-trained or non-co-trained policies111“Co-trained policies” refers to policies that have been trained together in the same environment with shared experiences. We refer the readers to Section 6.2 of Wang et al. (2023) for more information.. Experimental results on the StarCraft environment (Samvelyan et al. (2019)) demonstrate that PegMARL converges effectively even with demonstrations from non-co-trained policies.
2 Related Works
Imitation Learning (IL). IL methods seek to replicate an expert policy from demonstration data generated by that policy. Behavior Cloning (BC) is a commonly employed IL technique (Pomerleau (1991); Bojarski et al. (2016)), where the expert policy is estimated through supervised learning on demonstration data. However, BC is susceptible to compounding errors resulting from distribution shifts (Ross et al. (2011)). Another thread of IL research is Inverse Reinforcement Learning (IRL) (Ng et al. (2000); Ziebart et al. (2008)), in which the underlying reward function is estimated from the demonstration data and then used for policy learning. To alleviate the computational overhead of IRL, Generative Adversarial Imitation Learning (GAIL) (Ho & Ermon (2016)) was introduced, allowing direct policy learning from observed data without intermediate IRL steps. Song et al. (2018) extended this approach to introduce Multi-Agent GAIL (MAGAIL), which adapts GAIL to high-dimensional environments featuring multiple cooperative or competing agents. In general, IL approaches can rarely perform better than demonstrations. Therefore, they are not directly suitable for scenarios where only personalized expert demonstrations that do not demonstrate how to collaborate are available.
Learning from Demonstration (LfD). In contrast to IL methods, LfD aims to leverage demonstration data to facilitate learning rather than simply mimicking expert behavior. Existing LfD works incorporate demonstration data into a replay buffer with a prioritized replay mechanism to accelerate the learning process. Due to the off-policy nature of the demonstration data, most methods are value-based Hester et al. (2018); Vecerik et al. (2017). There has been some recent work where the demonstrations are used to aid exploration, especially in environments with large state-action spaces (Kang et al. (2018); Chen & Xu (2022); Rengarajan et al. (2022)). For instance, POfD (Kang et al. (2018)) learns an implicit reward from the demonstration data using a discriminator and incorporates it into the original sparse reward. LOGO (Rengarajan et al. (2022)) uses the demonstration data to directly guide the policy update: during each update iteration, the algorithm seeks a policy that closely resembles the behavior policy within a trust region. However, these approaches are primarily focused on single-agent settings.
In multi-agent settings, Qiu et al. (2022) suggests using demonstrations to pretrain agents through imitation learning as a warm start, followed by optimization of the pretrained policies using standard MARL algorithms. Lee & Lee (2019) augment the experience buffer with demonstration trajectories and gradually decrease the mixing of demonstration samples during training to prevent the learned policy from being overly influenced by demonstrations. Similar to POfD, DM2 (Wang et al. (2023)) enables agents to enhance their task-specific rewards by training discriminators as well. Each agent matches toward a target distribution of concurrently sampled trajectories from a joint expert policy to facilitate coordination. These approaches, however, require joint demonstrations of the same team configurations, which can be cumbersome (as we mentioned in section 1). Our approach differs in that we leverage only personalized expert demonstrations to learn a cooperative policy.
3 Preliminaries
Figure 2: An example of utilizing personalized demonstrations to learn cooperative multi-agent policies. To learn successful cooperation, the agents are required not only to imitate the demonstrations to achieve personal goals but also to learn how to avoid conflicts and collaborate. We visualize the state visitation frequency of the personalized demonstrations and the joint policies learned by our algorithm and MAPPO, where a darker color means a higher value. We observe that the demonstrations guide the agents in exploring the state space more efficiently than in MAPPO.
We start by considering a Markov Decision Process $(\mathcal{S},\mathcal{A},P,r,\gamma)$ for a cooperative multi-agent setting with $N$ agents. Here, $\mathcal{S}$ denotes the global state across all the agents, which can be decomposed as the product of $N$ local spaces $\mathcal{S}_{i}$ as $\mathcal{S}=\mathcal{S}_{1}\times\mathcal{S}_{2}\times\cdots\times\mathcal{S}_{N}$. By noting the local state of agent $i$ as $s_{i}\in\mathcal{S}_{i}$, the global state is $s=(s_{1},s_{2},\cdots,s_{N})$. Similarly, we define the global action space $\mathcal{A}$ as $\mathcal{A}=\mathcal{A}_{1}\times\mathcal{A}_{2}\times\cdots\times\mathcal{A}_{N}$, meaning that for any $a\in\mathcal{A}$, we may write $a=(a_{1},a_{2},\cdots,a_{N})$ with $a_{i}\in\mathcal{A}_{i}$. The transition probability from state $s$ to $s^{\prime}$ after taking a joint action $a$ is denoted by $P(s^{\prime}|s,a)=\prod_{i}P_{i}(s_{i}^{\prime}|s,a)$. All agents share a common reward function $r$, and $\gamma\in[0,1]$ is the discount factor. In this work, we focus on decentralized learning and define the global policy as $\pi_{\theta}(a|s)$, where $\theta\in\Theta$ are the policy parameters. Specifically, we have $\theta=(\theta_{1},\theta_{2},\cdots,\theta_{N})$ as the factorized global policy parameters, and we can write $\pi_{\theta}(a|s)=\prod_{i}\pi_{\theta_{i}}(a_{i}|s)$ using policy factorization.
Personalized Tasks and MDPs. To introduce the notion of personalized demonstration, we need to extract the individual tasks of each agent from the collective tasks of multiple agents. For example, in Figure 2, the green agent’s personalized task is to open the door and the red agent’s personalized task is to reach the key without the other’s presence. We then define Personalized Markov Decision Processes (PerMDPs) $(\mathcal{S}_{i},\mathcal{A}_{i},Q_{i},r_{i},\gamma)$ for each agent or, more generally, for each $type$ of agents that share the same objective. Here, a $type$ refers to behaviorally identical agents using the formalism from Bettini et al. (2023), and we provide the PerMDP definition for each agent for simplicity. We assume that the state space $\mathcal{S}_{i}$ and action space $\mathcal{A}_{i}$ of the personalized task $\mathcal{T}_{i}$ are the same as the local state and action spaces of agent $i$ from the joint task. The transition probability from state $s_{i}$ to $s_{i}^{\prime}$ after taking a action $a_{i}$ is represented as $Q_{i}(s^{\prime}_{i}|s_{i},a_{i})$. By following an arbitrary policy $\pi_{i}(a_{i}|s_{i})$ for this personalized task, we can collect a set of personalized demonstrations $\mathcal{B}_{i}=\{(s_{i}^{t},a_{i}^{t})\}_{t=0}^{H}$, where $H$ is the episode horizon. We’d like to emphasize that while we assume the personalized tasks and the joint task are conducted in the same environment map, the underlying transition dynamics are different. Additionally, the joint reward $r$ may not necessarily equal the summation of rewards $r_{i}$ for each personalized task.
Occupancy Measures. For a joint policy $\pi=(\pi_{1},\pi_{2},...,\pi_{N})$, we can write the global state action occupancy measure $\lambda^{\pi}(s,a)$ as
$\lambda^{\pi}(s,a)=\sum_{t=0}^{\infty}\gamma^{t}\cdot\mathbb{P}\Big{(}s^{t}=s,a^{t}=a\,\,\Big{|}\,\,\pi\Big{)}$ (1)
and write the corresponding local cumulative state-action occupancy measure as
$\lambda^{\pi}_{i}(s_{i},a_{i})=\sum_{t=0}^{\infty}\gamma^{t}\cdot\mathbb{P}\Big{(}s^{t}_{i}=s_{i},a^{t}_{i}=a_{i}\,\,\Big{|}\,\,\pi\Big{)}$ (2)
for $\forall a_{i}\in\mathcal{A}_{i},s_{i}\in\mathcal{S}_{i}$. An interesting observation to note here is that we can write the local occupancy measure as the marginalization of the global occupancy measure with respect to all other agents. Mathematically, it holds that
$\lambda^{\pi}_{i}(s_{i},a_{i})=\sum_{a\in\{a_{i}\}\times\mathcal{A}_{-i}}\sum_{s\in\{s_{i}\}\times\mathcal{S}_{-i}}\lambda^{\pi}(s,a)$ (3)
with $\mathcal{A}_{-i}\!\!=\!\!\Pi_{j\neq i}\mathcal{A}_{j}$ and $\mathcal{S}_{\!-i}\!\!=\!\!\Pi_{j\neq i}\mathcal{S}_{j}$.
4 MARL with Personalized Expert Demonstrations
Figure 3: When joint demonstrations are sampled from co-trained policies, the agents’ behaviors exhibit compatibility. In contrast, personalized demonstrations solely focus on how each agent achieves its individual goal and lack cooperative elements, potentially leading to conflicts.
Now, we are ready to present the main problem we are interested in solving in this work. Assume each agent $i$ is associated with a personalized task $\mathcal{T}_{i}$. We collect one set of expert demonstrations for each agent $i$ or, equivalently, each personalized task $\mathcal{T}_{i}$. By letting an expert user perform each personalized task in the respective personalized MDP, we obtain a collection of expert demonstrations denoted by $\{\mathcal{B}_{E_{1}},\mathcal{B}_{E_{2}},...,\mathcal{B}_{E_{N}}\}$. We assume that the underlying expert policy associated with $\mathcal{B}_{E_{i}}$ is $\pi_{E_{i}}(a_{i}|s_{i})$, and $\lambda^{\pi_{E_{i}}}$ is the occupancy measure following the expert’s policy $\pi_{E_{i}}$ for agent $i$.
Figure 4: A motivating example illustrating the imitation of personalized demonstrations in a multi-agent environment. The primary technical challenge lies in the discrepancy between the transition dynamics in the personalized MDP and the local transition dynamics for each agent in the multi-agent environment.
4.1 Formulating Personalized Expert-Guided MARL
In standard Multi-Agent Reinforcement Learning, agents aim to discover optimal joint policies that maximize the long-term return $R(\pi_{\theta})\mathrel{\mathop{\ordinarycolon}}=\langle\lambda^{\pi_{\theta}},r\rangle=\frac{1}{N}\sum_{i=1}^{N}\langle\lambda^{\pi_{\theta}}_{i},r\rangle$. Typically, the learning process commences with random exploration, which is often inefficient due to MARL’s exponentially growing exploration spaces, especially when rewards are sparse. Beginning with the intuition of leveraging personalized demonstrations as guidance for how each agent should accomplish their personalized tasks to promote more effective exploration, akin to Kang et al. (2018) for single-agent cases, we can define the objective for learning from personalized demonstrations in multi-agent settings as follows:
$\max_{\theta\in\Theta}\frac{1}{N}\sum_{i=1}^{N}\left(\langle\lambda^{\pi_{\theta}}_{i},r\rangle-\eta{{\mathbb{D}}_{\text{JS}}}\left(\lambda^{\pi_{\theta}}_{i}\leavevmode\nobreak\ ||\leavevmode\nobreak\ \lambda^{\pi_{E_{i}}}\right)\right),$ (4)
where $\eta$ is a weighting term balancing the long-term reward and the personalized policy similarity. The Jensen-Shannon (JS) Divergence terms enable individual agents to align their actions with their respective personalized demonstrations and facilitate the achievement of their specific objectives.
However, blindly imitating the personalized demonstration may not always yield favorable outcomes and can even impede the learning process. Previous works in MARL have predominantly utilized joint demonstrations as guidance. As demonstrated by DM2 (Wang et al. (2023)), trajectories must be sampled from a co-trained joint expert policy for their joint action-matching objective to converge. The crucial aspect that joint demonstrations offer, which personalized demonstrations lack, is compatibility. To illustrate, consider the example depicted in Figure 3, where two agents aim to swap positions without entering the danger region. Two sets of optimal joint policies exist for this task: agent $a$ takes path 1 and agent $b$ takes path 2, or vice versa. When sampling joint demonstrations from a set of co-trained policies, the agents’ behaviors will be naturally compatible. In contrast, if personalized demonstrations were provided, agents could indifferently choose between both paths to reach their goals, potentially resulting in conflicting strategies through naive imitation.
Hence, we pose the following question: How can we leverage the benefits of using personalized demonstrations to enhance exploration efficiency while circumventing the imitation of conflicting behaviors that might hinder joint-policy training?
Consider the scenario depicted in Figure 4, where the blue agent seeks to replicate its personalized demonstration within a multi-agent environment. The blue agent can successfully transition to the neighboring grid on the right if it’s unoccupied; otherwise, the transition fails. Viewing the blue agent as the focal agent and the green agent as part of the environment, successful imitation hinges on the agent’s “local transition dynamics” matching those of the environment where the personalized demonstrations originated. We define $P(s_{i}^{\prime}|s_{i},a_{i})$ as an approximation for the $i$-th agent’s local transition dynamics in the multi-agent environment, which is governed by the true transition dynamics $P(s^{\prime}|s,a)$ and the policy $\pi_{\theta}$. With this insight, we further refine the objective function as follows:
$\max_{\theta\in\Theta}\frac{1}{N}\sum_{i=1}^{N}\left(\langle\lambda^{\pi_{\theta}}_{i},r\rangle-\eta{{\mathbb{D}}_{\text{JS}}}\left(\hat{\lambda}^{\pi_{\theta}}_{i}\leavevmode\nobreak\ ||\leavevmode\nobreak\ \hat{\lambda}^{\pi_{E_{i}}}\right)\right),$ (5)
where $\hat{\lambda}^{\pi_{\theta}}_{i}$ and $\hat{\lambda}^{\pi_{E_{i}}}$ have the same definition as the original ${\lambda}^{\pi_{\theta}}_{i}$ and ${\lambda}^{\pi_{E_{i}}}$ but with a restricted domain:
$\text{dom}_{\hat{\lambda}^{\pi_{\theta}}_{i}}=\text{dom}_{\hat{\lambda}^{\pi_{E_{i}}}}=\{(s_{i},a_{i})\in\mathcal{S}_{i}\times\mathcal{A}_{i}|{\mathbb{D}}_{JS}\left(P_{i}(s_{i}^{\prime}|s_{i},a_{i})\leavevmode\nobreak\ ||\leavevmode\nobreak\ Q_{i}(s_{i}^{\prime}|s_{i},a_{i})\right)\leq\epsilon\}.$ (6)
where $Q(s_{i}^{\prime}|s_{i},a_{i})$ represents the true transition dynamics of the personalized MDP from which demonstrations are collected. By adjusting the learning objective in this manner, the occupancy matching only happens on those state-action pairs that guide us toward the desired next local state under the current policy.
4.2 Solving Personalized Expert-Guided MARL
Now, we can begin the development of a practical algorithm. By adopting Theorem 2 from Kang et al. (2018), we can substitute the JS divergence between occupancy measures with
${{\mathbb{D}}_{\text{JS}}}\left(\lambda^{\pi_{\theta}}_{i}\leavevmode\nobreak\ ||\leavevmode\nobreak\ \lambda^{\pi_{E_{i}}}\right)\approx\sup_{D_{i}}\left(\mathbb{E}_{{\pi_{\theta}}}[\log(1-D_{i}(s_{i},a_{i}))]+\mathbb{E}_{{\pi_{E_{i}}}}[\log(D_{i}(s_{i},a_{i}))]\right),$ (7)
where $D(s_{i},a_{i})\mathrel{\mathop{\ordinarycolon}}\mathcal{S}_{i}\times\mathcal{A}_{i}\rightarrow(0,1)$ is a discriminative classifier to discern if the $(s_{i},a_{i})$ pair is from the demonstration or the current policy. However, this doesn’t account for the constraint in Equation equation 6. Therefore, by introducing an indicator function $\mathds{1}(s_{i},a_{i})$ for whether $(s_{i},a_{i})$ belongs to the domain of $\hat{\lambda}^{\pi_{\theta}}_{i}$ and $\hat{\lambda}^{\pi_{E_{i}}}$, we can obtain
${{\mathbb{D}}_{\text{JS}}}\left(\hat{\lambda}^{\pi_{\theta}}_{i}\leavevmode\nobreak\ ||\leavevmode\nobreak\ \hat{\lambda}^{\pi_{E_{i}}}\right)\approx\sup_{D_{i}}\left(\mathbb{E}_{{\pi_{\theta}}}\left[\mathds{1}(s_{i},a_{i})\cdot\log(1-D_{i}(s_{i},a_{i}))\right]+\mathbb{E}_{{\pi_{E_{i}}}}\left[\mathds{1}(s_{i},a_{i})\cdot\log(D_{i}(s_{i},a_{i}))\right]\right).$ (8)
Since direct access to the transition distributions is unavailable, verifying whether each $(s_{i},a_{i})$ pair is within the domain is not directly feasible. To address this, we further approximate the indicator function using another discriminative classifier $D_{i}(s_{i},a_{i},s_{i}^{\prime})\mathrel{\mathop{\ordinarycolon}}\mathcal{S}_{i}\times\mathcal{A}_{i}\times\mathcal{S}_{i}\rightarrow(0,1)$, estimating the likelihood of a $(s_{i},a_{i},s_{i}^{\prime})$ tuple being from the demonstrations. In essence, it quantifies the likelihood that the corresponding $(s_{i},a_{i})$ induces a desired change in dynamics as observed in the demonstrations.
Algorithm 1 Personalized Expert-Guided MARL (PegMARL)
Input: Number of agents $N$; environment $env$; personalized expert trajectories $\mathcal{B}_{E_{1}},...,\mathcal{B}_{E_{N}}$; batch size $M$; weight parameters $\{\eta^{k}\}$.
Initialize: Policies $\{\pi_{\theta_{i}}\}$, discriminators $\{D_{\phi_{i}}\}$ and $\{D_{\bar{\phi}_{i}}\}$, where $i=1,2,...,N$.
Output: Learned policies $\{\pi_{\theta_{i}}\}$.
1:for iteration $k=0,1,2,...$ do
2:     Gather trajectories of multi-agent rollouts from $env$, $\mathcal{B}^{k}=ROLLOUT(\pi,env)$.
3:     for agent $i=0,1,...,N-1$ do
4:         Sample $M$ tuples of $(s_{i},a_{i},s_{i}^{\prime})$ from demonstration $\mathcal{B}^{E_{i}}$ and $\mathcal{B}^{k}$;
5:         Update personalized behavior discriminator $D_{\phi_{i}}$:
$\max_{\phi_{i}}\left(\mathbb{E}_{\mathcal{B}^{k}}\left[\log(1-D_{\phi_{i}}(s_{i},a_{i}))\right]+\mathbb{E}_{\mathcal{B}_{E_{i}}}\left[\log D_{\phi_{i}}(s_{i},a_{i})\right]\right).$ (9)
6:         Update personalized dynamics discriminator $D_{\bar{\phi_{i}}}$:
$\max_{\bar{\phi_{i}}}\left(\mathbb{E}_{\mathcal{B}^{k}}\left[\log(1-D_{\bar{\phi_{i}}}(s_{i},a_{i},s_{i}^{\prime}))\right]+\mathbb{E}_{\mathcal{B}_{E_{i}}}\left[\log D_{\bar{\phi}_{i}}(s_{i},a_{i},s_{i}^{\prime})\right]\right).$ (10)
7:         Estimate the reshaped reward as $\hat{r}_{i}^{k}=r-\eta^{k}D_{\bar{\phi_{i}}}(s_{i},a_{i},s_{i}^{\prime})\log(1-D_{\phi_{i}}(s_{i},a_{i}))$.
8:         Update agent policy $\pi_{\theta_{i}}$.
9:     end for
10:end for
After approximating the inner JS divergence term in Equation equation 5 with two discriminators, we integrate the discriminative rewards derived from these discriminators into the environmental reward for the outer problem. Specifically, we estimate the reshaped reward $\hat{r}_{i}$ as $\hat{r}_{i}=r-\eta D_{\bar{\phi_{i}}}(s_{i},a_{i},s_{i}^{\prime})\log(1-D_{\phi_{i}}(s_{i},a_{i}))$, where $\phi_{i}$ and $\bar{\phi_{i}}$ are parameters for the two discriminators. The personalized behavior discriminator $D_{\phi_{i}}$ evaluates local state-action pairs, providing positive incentives for actions that align with the demonstration and negative incentives for divergent ones, while the personalized dynamics discriminator $D_{\bar{\phi_{i}}}$ assesses if a local state-action pair induces a desired change in dynamics akin to that observed in the demonstration, adjusting the incentive weight accordingly. Subsequently, policy optimization is conducted by maximizing the long-term return with the reshaped reward. Our PegMARL algorithm, detailed in Algorithm 1, is compatible with any policy gradient methods, with MAPPO (Yu et al. (2021)) adopted in our implementation.
5 Experiments
In this section, we empirically evaluate the performance of PegMARL, focusing on the following questions: (1) How does PegMARL, which leverages personalized demonstrations, compare to state-of-the-art MARL techniques? (2) How does PegMARL scale with an increasing number of agents and in the case of continuous state-action spaces? (3) How does the sub-optimality of personalized expert demonstrations affect the performance of PegMARL? and (4) How does PegMARL perform when trained with joint demonstrations sampled from co-trained or non-co-trained expert policies?
5.1 Main Results with Personalized Demonstrations
We evaluate PegMARL with personalized demonstrations on both discrete gridworld environments (Figure 5) and a continuous multi-agent particle environment (Figure 7(a)). Our comparison includes four strong baselines: MAPPO (Yu et al. (2021)), a leading decentralized MARL algorithm; MAGAIL (Song et al. (2018)), a state-of-the-art multi-agent imitation learning algorithm; DM2 (Wang et al. (2023)), which combines distribution matching reward with environmental reward; and ATA (She et al. (2022)), one of the best multi-agent reward-shaping methods. Notably, MAGAIL and DM2 were not originally designed for personalized demonstrations. Hence, we have adapted them for use in personalized demonstration settings through necessary modifications.
(a) The lava scenario: the agents are homogeneous, aiming to reach corresponding diagonal positions without entering the lava.
Refer to caption Refer to caption Refer to caption
2 agents 3 agents 4 agents
Refer to caption Refer to caption
easy hard
(a) The lava scenario: the agents are homogeneous, aiming to reach corresponding diagonal positions without entering the lava.
(b) The door scenario: the agents are heterogeneous, the assistant agent (green) must reach the green square and remain there to hold the door open, allowing the other agent (red) to reach the goal.
Figure 5: The discrete gridworld environments. Circles indicate the starting locations of the agents, while squares of the same color denote their respective goal locations. More details can be found in Appendix A.1.
Figure 6: Learning curves of PegMARL versus other baseline methods under the lava scenario. PegMARL converges to higher rewards and generalizes better to larger numbers of agents. The star symbols (*) in the legend indicate that suboptimal personalized demonstrations are adopted.
Figure 7: Learning curves of PegMARL versus other baseline methods under the door scenario. PegMARL shows better robustness in terms of convergence and generalizability.
For each scenario, we collect two sets of personalized demonstrations for each agent: optimal and suboptimal. The average episodic rewards of suboptimal demonstrations are approximately half of their optimal counterparts (more details can be found in Appendix B). We execute each method in every environment with 10 distinct random initializations and plot the mean and variance across all runs. Oracle denotes the best possible return achievable with optimal policies.
How does PegMARL scale with an increasing number of agents? The lava scenario (Figure 5) contains three variations, each involving different quantities of agents and escalating levels of complexity. The agents receive a positive reward only upon reaching the goal, incur penalties for collisions, and instantly die and terminate the episodes if they step into the lava, making learning in this environment challenging. As we observe from Figure 4(b), MAPPO struggles to develop meaningful behavior across all scenarios due to the sparse reward structure. MAGAIL performs worse, primarily due to the absence of environmental reward signals. The naive imitation of personalized agent demonstrations leads to frequent collisions among agents. While ATA can learn suboptimal policies in the 2-agent setting, its learning efficacy declines drastically as the number of agents increases. This underscores the challenges of training as the complexity of the multi-agent environment increases.
In contrast, PegMARL demonstrates superior generalizability across scenarios with more agents, maintaining stable performance despite suboptimal demonstrations. As DM2 has been adapted to accommodate personalized demonstrations, it closely resembles PegMARL but lacks the personalized dynamics discriminator component. As the number of agents increases, we would expect the number of ways for agents to interact to increase, and therefore for the personal demonstrations to be potentially misleading. Consequently, DM2 experiences a decrease in convergence speed as the number of agents increases. This highlights the importance of the personalized dynamics discriminator in PegMARL, enabling effective handling of inter-agent interactions in personalized demonstration scenarios, thereby ensuring its efficacy across varying agent counts.
How does PegMARL perform under the heterogeneous setting? The door scenario (Figure 5) contains two variants with varying levels of difficulty. The easy case is fairly straightforward: success should be achievable by adhering to personalized demonstrations, which illustrate how each agent navigates to their respective goal locations. As shown in Figure 7, most algorithms, except for MAGAIL, showcase proficient performance. Notably, PegMARL exhibits the swiftest convergence. We attribute MAGAIL’s failure in this case to its inability to direct the green agent to remain positioned at the green square — a behavior not explicitly demonstrated. This underscores the importance of environmental reward signals when integrating personalized demonstrations into multi-agent learning paradigms. The hard case necessitates a higher degree of agent cooperation: once the red agent gains entry to the middle room, the green agent must move aside from the green square to enable the red agent’s passage into the right room. In this complex setting, only PegMARL and DM2 demonstrate commendable convergence. PegMARL, equipped with the personalized dynamics discriminator, maintains faster convergence compared to DM2.
How does PegMARL perform in the continuous setting? We modified the cooperative navigation task from the multi-agent particle environment (Lowe et al. (2017); Mordatch & Abbeel (2017)) to evaluate the performance of our algorithm in a continuous environment (more details can be found in Appendix A.2). Figure 7(b) demonstrates the learning curves of PegMARL versus MAPPO and DM2. This validates the ability of PegMARL in continuous environments.
(a) The modified cooperative navigation scenario: the agents need to navigate around the wall to occupy both landmarks.
(b) Learning curves of PegMARL versus MAPPO and DM2 under the modified cooperative navigation scenario
Figure 8: (a) The modified cooperative navigation scenario. (b) The learning curves demonstrate that PegMARL is effective in continuous environments.
5.2 Results with Joint Demonstrations
While PegMARL is primarily designed to leverage personalized demonstrations, it can also be extended to utilize joint demonstrations. Unlike other MARL algorithms like DM2 (Wang et al. (2023)) that require compatible joint demonstrations sampled from co-trained policies to achieve convergence, PegMARL has the flexibility to utilize joint demonstrations even from non-co-trained policies. To demonstrate, we conduct comparative experiments within the StarCraft Multi-Agent Challenge (SMAC) environment (Samvelyan et al. (2019)). We employed the same two tasks that were used in DM2: 5mv6m, which involves 5 Marines (allies) against 6 Marines (enemies), and 3sv4z, which features 3 Stalkers (allies) versus 4 Zealots (enemies). The win rates of the demonstrations in both cases are approximately $30\%$.
Figure 9: Learning curves of PegMARL versus DM2 in two tasks under the SMAC scenarios. The suffix “diff” in the legend indicates that the joint demonstrations used are sampled from non-co-trained policies. Otherwise, the demonstrations are sampled from co-trained policies.
How does PegMARL scale with joint demonstrations? Figure 9 presents the learning curves of PegMARL alongside the DM2 baseline in both the 5mv6m and 3sv4z configurations within the SMAC environment. When the joint demonstrations are sampled from co-trained policies, PegMARL achieves a comparable, and in some cases even higher, success rate compared to DM2 in both tasks.
When the joint demonstrations are sampled from policies that are non-co-trained, the success rate of DM2 exhibits a significant decline in both tasks. In the 5mv6m task, DM2’s success rate drops to nearly 0, indicating a failure to learn. In contrast, PegMARL maintains a similar level of performance compared to results from co-trained demonstrations, with only a slight decrease in convergence speed. These findings underscore the versatility and effectiveness of PegMARL in scenarios requiring collaboration among diverse agents, demonstrating its robustness across various multi-agent environments.
6 Conclusions
In this work, we introduce PegMARL, a novel approach for Multi-Agent Reinforcement Learning, which adopts personalized expert demonstrations as guidance and allows agents to learn to cooperate. Specifically, the algorithm utilizes two discriminators to dynamically reshape the reward function: one provides incentives to encourage the alignment between policy behavior with provided demonstrations, and the other regulates incentives based on whether the behavior leads to the desired objective. We demonstrate PegMARL’s effectiveness, scalability, and robustness in both discrete and continuous environments, where it outperforms state-of-the-art decentralized MARL algorithms, pure imitation learning, and reward-shaping techniques. We observe that PegMARL can achieve near-optimal policies even with suboptimal demonstrations. Furthermore, We showcase PegMARL’s capability to leverage joint demonstrations and converge successfully, regardless of whether they are sampled from co-trained or non-co-trained policies.
Appendix A Environment Details
A.1 Discrete Gridworld Environments
We use two gridworld environments with discrete state and action space: the lava scenario and the door scenario (Figure 5). The agents in the lava scenario are homogeneous because they have the same objective: navigating to their corresponding goal location. The door scenario is heterogeneous: the assistant agent (green) must open the door while the other agent (red) must reach the goal.
  • The lava scenario: This environment has a 6-by-6 lava pond in the center (Figure 5). We provide three variations of this scenario, each with varying numbers of agents and increasing levels of complexity. The main goal for the agents is to efficiently navigate to their assigned goal locations while avoiding stepping into the lava. An episode terminates when all agents reach their respective goals (succeed), or if any agents step into the lava or the maximum episode length is reached (fail).
  • The door scenario: This environment is adapted from Franzmeyer et al. (2022) (see Figure 5). In this scenario, the green agent must navigate to a designated green square and maintain its presence there to sustain the open state of the green door, thereby enabling the entry of a red agent into the right side room. An episode ends when the red agent reaches the red goal location (succeed) or the maximum episode length is reached (fail).
[…]
A.2 Cooperative Navigation
We modified the cooperative navigation task of multi-agent particle environment Lowe et al. (2017); Mordatch & Abbeel (2017) to evaluate the performance of our algorithm in a continuous environment. The modified environment consists of 2 agents, 2 goal landmarks and 1 wall between the agents and the goals. The agents need to navigate around the wall to occupy both landmarks. The state definition is the same as the original. We sparsify their original reward as follows:
$r=\sum_{a\in A}\min(0.3,\min_{g\in G}d(a,g))+\sum_{g\in G}\min(0.3,\min_{a\in A}d(a,g)).$ (11)
Appendix B Additional Experiments Details and Visualizations
Figure 10 shows an example of personalized demonstrations for the lava scenario. Table 1 summarizes the details of the suboptimal demonstrations, whose average episodic rewards are around 4.5 and about half of their optimal counterparts.
[Uncaptioned image] Figure 10: An example of personalized demonstrations for the lava scenario (we did not visualize all the optimal paths). There is only one agent in the environment. The agent may take both the left and the right path toward the goal.
Agent Id $\mathcal{S}$ $\mathcal{A}$ Samples Average Episodic Reward 1 $\mathbb{R}^{2}$ 5 300 4.42 2 $\mathbb{R}^{2}$ 5 440 4.79 3 $\mathbb{R}^{2}$ 5 344 4.3 4 $\mathbb{R}^{2}$ 5 360 4.41 (a) The lava scenario Case Agent $\mathcal{S}$ $\mathcal{A}$ Samples Average Episodic Reward easy red $\mathbb{R}^{2}$ 5 643 4.01 green $\mathbb{R}^{2}$ 5 612 4.39 hard red $\mathbb{R}^{2}$ 5 607 4.34 green $\mathbb{R}^{2}$ 5 593 4.36 (b) The door scenario
Table 1: The details of suboptimal demonstrations.
We provide visual representations of the policy and state occupancy measure corresponding to suboptimal demonstrations in Figure 11 for the lava scenarios and in Figure 12 for the door scenario. The red square symbolizes the agent’s initial position in these visualizations, while the green square designates its respective goal location. Arrows within the figures denote available actions at each state, with arrow length indicating the probability associated with each action.
(a) Agent 1
(b) Agent 2
(c) Agent 3
(d) Agent 4
Figure 11: Agent policies and state occupancy measures estimated from the suboptimal demonstrations for the lava scenario.
(a) The easy case
(b) The hard case
Figure 12: Agent policies and state occupancy measures estimated from the suboptimal demonstrations for the door scenario. The top row is for the red agent, and the bottom row is for the green agent.
We additionally depict the state visitation frequencies of the joint policies learned by PegMARL with suboptimal demonstrations and MAPPO for both the lava (Figure 13) and the door scenario (Figure 14).
(a) 2-agent case
(b) 3-agent case
Refer to caption Refer to caption
PegMARL MAPPO
Refer to caption Refer to caption
PegMARL MAPPO
Refer to caption Refer to caption
PegMARL MAPPO
(a) 2-agent case
(b) 3-agent case
(c) 4-agent case
Figure 13: State visitation frequency of the joint policies learned by PegMARL (with suboptimal demonstrations) and MAPPO for the lava scenario. The darker color means a higher value. MAPPO failed to learn any meaningful policies in all three settings.
(a) The easy case
Refer to caption Refer to caption
PegMARL MAPPO
Refer to caption Refer to caption
PegMARL MAPPO
(a) The easy case
(b) The hard case
Figure 14: State visitation frequency of the joint policies learned by PegMARL (with suboptimal demonstrations) and MAPPO for the door scenario. The darker color means a higher value. MAPPO converges to a suboptimal policy in the easy case, while failing to learn in the hard case.